# MIST: A Simple and Scalable End-To-End 3D Medical Imaging Segmentation Framework

### v1

@@ 1 Introduction
To address this inconsistency, we propose the Medical Imaging Segmentation Toolkit or MIST, a simple, modular, and end-to-end deep learning-based medical imaging segmentation framework that allows researchers to seamlessly train, test, and evaluate existing and new methods on their data in a consistent and reproducible way. MIST has a rule-based data analysis and preprocessing pipeline that it makes available to multiple architectures, loss functions, and training parameters. Additionally, MIST has a standardized and flexible evaluation pipeline that computes multiple metrics like the Dice, Haudorff distance, and surface Dice for any user-defined segmentation class. This standardization of several pipelines allows for fair comparisons between methods and their effect on segmentation accuracy for any given dataset.
@@ 2.1 Data Format
MIST uses the same file structure as the BraTS challenge datasets [28]. That is, MIST assumes that the train and test (optional) data directories are formatted so that each sub-directory corresponds to one patient and that each imaging modality (i.e., T1, T2, or CT) is its own NIfTI file. The labels should all be in a single NIfTI file in the same sub-directory for each patient. Once the dataset is in the correct format, the next step is to prepare a small JSON file containing the details of the dataset. These include information like the imaging modality, naming conventions for the images and masks, labels, and the final classes that we use for evaluation (i.e., labels 1, 2, and 3 for whole tumor). An example of this JSON file is provided in the documentation page here. MIST also provides support for converting datasets that are in CSV or MSD format.
@@ 2.2.1 Data Analysis
Often, zeros or useless background information surround the region of interest. For example, zeros surround the region outside of the brain in the BraTS datasets, and the table and air surrounding the body in CT imaging do not provide relevant information. From a computational perspective, it is desirable to crop the images to the foreground to make them smaller. Additionally, cropping to the foreground also reduces the sparsity of the dataset. MIST crops each image by first windowing it with its 33rd and 99.5 percentile values, then using an Otsu filter to produce a foreground threshold. We then crop the image according to the binary mask resulting from this threshold (i.e., one if the voxel intensity is greater or equal to the threshold). To determine if we need to use cropping, we check to see if, on average, the resulting volume reduction is at least 20%.
@@
To determine the target spacing for a dataset, we first look at the distribution of the voxel spacings along each axis. Our initial target spacing is the median spacing in each direction. However, suppose the resulting target spacing is anisotropic (i.e., the maximum and minimum spacing ratio is greater than three), then we take the 10th percentile spacing along the lowest resolution axis in that case. However, if resampling with this selected target spacing results in any image taking up more than 2 GB of memory, then we increase all components of our target spacing by 25%. We repeat this process until none of the resampled images take up more than 2 GB of memory. This last step is important because it conserves computational resources. Larger images take longer to load from disk and require significantly more resources to manipulate (i.e., patch extraction and augmentation).
@@
In addition to the previous analysis steps, MIST checks to see that the header information in the images and masks agree and, if there are multiple modalities, that the headers in each image agree. If there is a disagreement between the headers, MIST prints a warning with the patient ID to the console and excludes the example from training. MIST also computes equal-weighted class weights for each class based on the amount of voxels belonging to each class over the entire dataset. These class weights are primarily informational but can be used for weighted loss functions.
@@ 2.2.3 Training
Once preprocessing is complete, MIST uses a five-fold cross-validation to train and evaluate a model on the dataset. In other words, for each fold we set aside 20% of the data as an independent test set and use the other 80% of the data for training. By default, MIST sets aside 10% of the training data as a validation set, but users can adjust this percentage. Additionally, users have the option to adjust the number of folds and can even specify custom folds if, for example, they are running leave-one-institution-out cross-validation.
@@
MIST supports six different network architectures. These architectures include nnUNet (default) [10], U-Net [23, 7], Swin UNETR [2], and PocketNet [3]. MIST also provides users with options like adding deep supervision [30] or variational autoencoder regularization [25] to most of these architectures. Other regularization features like L2 and L1 regularization and gradient clipping are available for all architectures. For more details about the supported architectures (and their differences), please refer to our documentation page. Finally, MIST’s modular design allows users to implement their own architectures and integrate them into MIST with minimal effort.
@@
Like with the choice of network architecture, MIST allows users to choose from a variety of loss functions like Dice with Cross Entropy (default) [22], the Generalized Dice Loss [24], and boundary-based loss functions like the Boundary, Hausdorff, and Generalized Surface losses [14, 13, 5]. Again, MIST’s modular design allows researchers to implement custom loss functions and seamlessly use them within the MIST framework.
@@ 2.2.4 Auxiliary Pipelines
Once training is complete, the evaluation pipeline launches to assess the accuracy of the predictions with respect to the final classes defined by the user in their dataset JSON file. By default, these metrics are the Dice similarity coefficient and the 95th percentile Hausdorff distance (HD95). Other metrics like the surface Dice and average surface distance are also available. The output of the evaluation pipeline is a CSV file called results.csv where each row shows the accuracy for each class and metric for one patient. The last five rows give the mean, standard deviation, median, and 25th and 75th percentiles for all of the metrics and classes. The evaluation pipeline is also available as a stand-alone pipeline that can run on any set of predictions.
@@
Postprocessing is available as an optional stand-alone pipeline. This pipeline allows users to specify which labels to apply a user-specified set of operations. These operations include the removal of small objects, only keeping the top-$k$ components (adjustable value of $k$), morphological cleaning, and filling holes with a specified value. Once the postprocessing pipeline finishes, the evaluation pipeline launches and compares the change in accuracy to the baseline results. MIST computes this change in accuracy as a weighted average of the specified metrics.

### v2

@@ 1 Introduction
To address this inconsistency, we propose the Medical Imaging Segmentation Toolkit or MIST, a simple, modular, and end-to-end deep learning-based medical imaging segmentation framework that allows researchers to seamlessly train, test, and evaluate existing and new methods on their data in a consistent and reproducible way. MIST has a rule-based data analysis and preprocessing pipeline that it makes available to multiple architectures, loss functions, and training parameters. Additionally, MIST has a standardized and flexible evaluation pipeline that computes multiple metrics like the Dice, Hausdorff distance, and surface Dice for any user-defined segmentation class. This standardization of several pipelines allows for fair comparisons between methods and their effect on segmentation accuracy for any given dataset.
@@ 2.1 Data Format
MIST uses the same file structure as the BraTS challenge datasets [35, 15]. That is, MIST assumes that the train and test (optional) data directories are formatted so that each sub-directory corresponds to one patient and that each imaging modality (i.e., T1, T2, or CT) is its own NIfTI file. The labels should all be in a single NIfTI file in the same sub-directory for each patient. Once the dataset is in the correct format, the next step is to prepare a small JSON file containing the details of the dataset. These include information like the imaging modality, naming conventions for the images and masks, labels, and the final classes that we use for evaluation (i.e., labels 1, 2, and 3 for whole tumor). An example of this JSON file is provided in the documentation page here. MIST also provides support for converting datasets that are in CSV or MSD format.
@@ 2.2.1 Data Analysis
Often, zeros or useless background information surround the region of interest. For example, zeros surround the region outside of the brain in the BraTS datasets, and the table and air surrounding the body in CT imaging do not provide relevant information. From a computational perspective, it is desirable to crop the images to the foreground to make them smaller. Additionally, cropping to the foreground also reduces the sparsity of the dataset. MIST crops each image by first windowing it with its 33rd and 99.5 percentile values, then using an Otsu filter to produce a foreground threshold. Note that the windowing percentiles were determined experimentally with several datasets like LiTS, MSD, NFBS, BraTS, and AMOS [2, 28, 14]. We then crop the image according to the binary mask resulting from this threshold (i.e., one if the voxel intensity is greater or equal to the threshold). To determine if we need to use cropping, we check to see if, on average, the resulting volume reduction is at least 20%.
@@
To determine the target spacing for a dataset, we first look at the distribution of the voxel spacings along each axis. Our initial target spacing is the median spacing in each direction. However, suppose the resulting target spacing is anisotropic (i.e., the maximum and minimum spacing ratio is greater than three), then we take the 10th percentile spacing along the lowest resolution axis in that case. However, if resampling with this selected target spacing results in an example taking up more than 2 GB of memory, then MIST will print a warning to users asking them to consider coarsening the image. We set this 2 GB threshold to conserve computational resources. Larger images take longer to load from disk and require more resources to manipulate (i.e., patch extraction and augmentation).
@@
In addition to the previous analysis steps, MIST checks to see that the header information in the images and masks agree and, if there are multiple modalities, that the headers in each image agree. If there is a disagreement between the headers, MIST prints a warning with the patient ID to the console and excludes the example from training.
@@ 2.2.3 Training
Once preprocessing is complete, MIST uses a five-fold cross-validation to train and evaluate a model on the dataset. In other words, for each fold MIST sets aside 20% of the data as a validation dataset and uses the remaining 80% of the data for training. Users can adjust these percentages, specify a different number of folds, or even specify custom folds if, for example, they are running leave-one-institution-out cross-validation.
@@
MIST supports six different network architectures. These architectures include nnUNet (default) [11], U-Net [29, 8], Swin UNETR [3], PocketNet [4], and MedNeXt [30]. MIST also provides users with options like adding deep supervision [37] or variational autoencoder regularization [32] to most of these architectures. Other regularization features like L2 and L1 regularization and gradient clipping are available for all architectures. For more details about the supported architectures (and their differences), please refer to our documentation page. Finally, MIST’s modular design allows users to implement their own architectures and integrate them into MIST with minimal effort.
@@
Like with the choice of network architecture, MIST allows users to choose from a variety of loss functions like Dice with Cross Entropy (default) [26], the clDice loss [31], and boundary-based loss functions like the Boundary, Hausdorff, and Generalized Surface losses [17, 16, 6]. Again, MIST’s modular design allows researchers to implement custom loss functions and seamlessly use them within the MIST framework.
@@ 2.2.4 Auxiliary Pipelines
Once training is complete, the evaluation pipeline launches to assess the accuracy of the predictions with respect to the final classes defined by the user in their dataset JSON file. By default, these metrics are the Dice similarity coefficient and the 95th percentile Hausdorff distance (HD95). Other metrics like the surface Dice and average surface distance are also available. The output of the evaluation pipeline is a CSV file called results.csv where each row shows the accuracy for each class and metric for one patient. The last five rows give the mean, standard deviation, median, and 25th and 75th percentiles for all of the metrics and classes. The evaluation pipeline is also available as a stand-alone pipeline that can run on any set of predictions. With the exception of the surface Dice, all of our metrics are in line with Metrics Reloaded [25].
@@
Postprocessing is available as an optional stand-alone pipeline. This pipeline allows users to specify which labels to apply a user-specified set of operations. These operations include the removal of small objects, only keeping the top-$k$ components (adjustable value of $k$), morphological cleaning, and filling holes with a specified value. Once the postprocessing pipeline finishes, the evaluation pipeline launches and compares the change in accuracy to the baseline results. MIST computes this change in accuracy as a weighted average of the specified metrics. Note that MIST applies postprocessing to the discrete predictions, not the softmax probabilities.
@@ 4 Discussion
Future work will include an ablation study of the parameters that we selected for the BraTS dataset, like the use of deep supervision, residual convolution blocks, L2 regularization, and pocket architectures. Additionally, an in-depth comparison between other frameworks like nnUNet, MedNeXt, and Swin UNETR for this dataset will be an objective of future work. For an in-depth comparison of these frameworks for a liver segmentation dataset, please refer to [27].